# Contactless Fingerprint Recognition Using 3D Graph Matching

### v1

@@ IV-B Evaluation of Feature Network
All
4.9308 1.0821 0.3004 41.3399
Front Pose
4.4374 1.0109 0.2943 44.2322
Side Pose
5.1831 1.1886 0.3036 39.8613
CFPose
All
5.8304 1.4767 0.2909 34.7563
Front Pose
5.7583 1.4631 0.2774 30.6444
Side Pose
5.9965 1.5078 0.3221 44.2271
Table IV shows the minutiae extraction accuracy. As in [11], we use the precision, recall, and F1 score to examine the minutiae accuracy. A true positive detection is defined as the distance between the output minutia location and ground-truth minutia location is less than 8 pixels, and the angle difference is less than $15^{\circ}$. Otherwise, the detection would be false positive.
It should be noted that the ground truth depth and gradient maps are generated using the shape-from-silhouette method in [18]. Therefore, they are theoretically synthetic data. As for orientation, ridge period, and minutiae map, their ground-truth values are extracted by VeriFinger without manual labeling. Therefore, they may not be real values. Evaluation of these features is just a references here to exam whether the network has produced reasonable features for the matching algorithm.
Especially, the VeriFinger is originally designed for contact-based fingerprints, and fails on extracting correct minutiae on some of the contactless fingerprints. Therefore, as Table IV shows, the minutiae extraction accuracy comparing to VeriFinger is not very high, as the VeriFinger itself may not extract correct minutiae.
TABLE IV: Evaluation results of minutiae extraction on the UWA and CFPose database
Database Poses Precision Recall F1-score
UWA
All
53.25% 62.21% 56.60%
Front Pose
58.24% 63.26% 60.07%
Side Pose
50.70% 61.68% 54.83%
CFPose
All
47.75% 85.34% 60.65%
Front Pose
48.67% 85.33% 61.37%
Side Pose
45.62% 86.01% 58.99%
IV-C Evaluation of Matching Network
IV-C1 Matching Protocols
Matching experiments are performed on UWA [41] and CFPose [11]. Contactless fingerprint matchings are performed, as the proposed method aims at utilizing 3D features to improve contactless fingerprint matching performances, which is in contrast to many previous methods that aim at contactless-contact matching. These CL2C methods also do not solve the pose variation issue in CL2CL matching.
UWA Database. Matching experiments are performed on the first 5,970 fingerprints from 1,500 fingers, of which almost every finger has 3 poses $\times$ 2 images. For the 5,970 images, due to the different matching protocols of the current two best methods [18][17], we conduct experiments on this dataset separately according to these two protocols. Following [18], 3,000$\times$3,000 matchings are performed. It should be noticed that matchings of different poses from the same finger are also genuine matching. There are 9,000 pairs of genuine matchings, and the rest of the pairs are impostor matching.
As for [17], Dong et al. only utilized the 2,988 images from the first capture, which included 3 poses $\times$ 1 image. In our experiments, we consider the number of images to be 3,000. We conduct identification performance experiments with totaling 1,000$\times$2,000 matching attempts, and matching performance experiments with totaling 3,000$\times$2,999///2 matching attempts. For the missing data in the dataset, we set their matching scores to 0.
CFPose Database. This database contains 1,400 fingerprints from 140 fingers, and each finger has 10 images. We use the first 1,200 images for matching. For evaluating the matching performance, a total of 719,400 pairs of matchings are performed. 5,400 pairs are genuine matching, and the rest are impostor matching. For evaluating the identification performance like Dong et al. [17], a total number of 120 $\times$ 1080 $=$ 129,600 pairs of matchings are performed. 1,080 pairs are genuine matching, and the rest are impostor matching.
IV-C2 Matching Results
TABLE V: Matching results on the UWA database
Method EER Rank-1
Verifinger 21.19%1 97.13%1
37.73%2 65.60%2
Tan et al. [11]
14.27%1† 87.29%1†
Cui et al. [18] 14.12%1† 96.70%1
24.27%2 88.00%2
Dong et al. [17]
13.84%1† 56.07%2†
Proposed 8.66%1 96.73%1
9.16%2 88.35%2
[…]
  • 1: We sample 3 poses for the first 1,000 fingerprint IDs, each with 2 images, and conduct 3,000$\times$3,000 matchings and identification experiments following the test protocol in [18].
  • 2: We sample 3 poses for the first 1,000 fingerprint IDs, each with 1 image, and conduct 3,000$\times$2,999///2 matchings and 1,000$\times$2,000 identification experiments following the test protocol in [17]. In the identification experiment, impressions with the least pose variation are served as the gallery and those remained are the probe. For the missing 4 fingerprint IDs, we set their matching scores with any other fingerprint to 0.
TABLE VI: Matching results on the CFPose database
Method EER Rank-1
Verifinger 7.13%1 100.00%1
93.06%2
Tan et al. [11] 1.95%1† 99.35%1†
99.17%2
Dong et al. [17]
1.12%1† 96.70%2†
Proposed 1.74%1 99.92%1
99.17%2
[…]
  • 1: We sample 10 poses for the first 120 fingerprint IDs, and conduct 1,200$\times$1,199///2 matchings and identification experiments following the test protocol in [11].
  • 2: We sample 10 poses for the first 120 fingerprint IDs, and conduct 120$\times$1,080 identification experiments following the test protocol in [17]. Impressions with the least pose variation are served as the gallery and those remained are the probe.
We compare our matching results with the current SOTA contactless-concactless fingerprint matching algorithms, which are shown in Table V and Table VI due to the different test protocols in [18] and [17]. The data marked with $\dagger$ in the two tables are the results obtained by reproducing previous methods using the latest version of Verifinger during our experiments. Due to the version upgrade of Verifinger and differences in parameter settings during matching, these results may differ from those in the reference literature. For example, in Table V, the results from Verifinger (37.73) show a much higher EER than the 26.38 reported by Dong et al. [17]. Still, the Rank-1 accuracy (65.60) is also significantly higher than their reported 50.05.
The experimental results show that our algorithm has achieved state-of-the-art on the UWA database and competitive results on the CFPose dataset. Table V demonstrates that our method, compared to the previous best method, achieves a decrease of 5.82% in EER for the UWA dataset, while the Rank-1 accuracy remains almost unchanged. For the CFPose dataset, our method reduces the EER by 0.21%, with the Rank-1 accuracy staying largely the same.
However, the method of Dong et al. [17] faces more incredible difficulty during testing, as there are no fingerprints with the same direction in the UWA dataset, which increases the difficulty of matching and identification. In this scenario, our method achieves a 5.18% reduction in EER and a 0.30% increase in Rank-1 accuracy compared to the previous best method on the UWA dataset. Moreover, the Rank-1 accuracy on the CFPose dataset also improves. Our method’s performance in terms of EER on the CFPose dataset is slightly inferior to that of the previous best method, and this may be because the previous method learned more discriminative information from fingerprints with smaller pose spans. In contrast, a higher proportion of fingerprints in our training set have larger pose spans, resulting in less discriminative information learned from fingerprints with smaller pose spans.
It can be observed that most of the time, our method achieves a slight improvement in Rank-1 accuracy while significantly reducing EER. This improvement indicates that our approach can better classify contactless fingerprints belonging to the same finger but with different directions compared to other methods. Mainly, it performs well in fingerprints with a large range of poses.
Our matching and identification performance on the UWA and CFPose datasets are illustrated in Fig. 9, 10, and 11, respectively. These figures compare our results before and after fusion with Verifinger. It is evident from the figures that our method demonstrates a significant improvement in reducing EER.
IV-C3 Ablation Studies
We conduct ablation studies on different training and fusion methods, and compare 3D minutiae to plain 2D minutiae.
Different training strategies. We conduct ablation experiments on the UWA dataset with different training strategies mentioned in this paper, as shown in Table VII and Fig. 12. It can be seen that when only fingers in a unified pose are used for training, the network finds it challenging to recognize that fingerprints in different poses belong to the same finger. Training fingerprints in various poses and fingerprints in the same pose in the original dataset ratio will make the network unable to converge effectively. After increasing the proportion of fingerprints in the same pose, the network can converge effectively, and the EER drops significantly, but the Rank-1 accuracy also drops significantly. Using 3D affine transformation to correct fingerprint poses and using the feature extraction network proposed in this paper can effectively reduce the EER, but slightly reduces the Rank-1 accuracy.
However, a simple score fusion method can combine the advantages of the two training methods to achieve good experimental results. It is worth noting that the score fusion method can even achieve a higher Rank-1 accuracy than the original network trained in the same pose, and the loss of EER is also tiny. The fusion results at FMR $=10^{-2}$ in the DET curve are poor, indicating that the score fusion algorithm is too simple and there is still room for improvement.
TABLE VII: Ablation study of training strategies on the UWA database
a b c d e f EER Rank-1
33.79% 55.79%
14.79% 56.07%
13.92% 42.95%
12.70% 40.99%
12.96% 57.70%
  • a: Using minutiae extracted by Verifinger.
  • b: Using minutiae extracted by Feature Network.
  • c: Training with same pose data.
  • d: Training with data equalization.
@@ IV-C3 Ablation Studies
  • f: Score fusion.
Comparison of 3D and 2D minutiae. To validate the superiority of 3D minutiae over 2D minutiae in CL2CL matching, we train a graph embedding model for only 2D minutiae under similar conditions: we used the same dataset for pretraining and finetuning, but during training, each minutia only included 2D position coordinates $\mathbf{p}^{\prime}=(x^{\prime},y^{\prime})$ and orientation $\theta^{\prime}$. For the spatial transformation network during finetuning, we only predicted 2D displacements $\mathbf{t}^{\prime}=(x_{t}^{\prime},y_{t}^{\prime})$ and rotation $\theta_{t}^{\prime}$. The obtained corrected 2D minutiae would be:
$\mathbf{m}_{t}^{\prime}=\begin{pmatrix}-\cos{\theta^{\prime}}&\sin{\theta^{\prime}}\\ \sin{\theta^{\prime}}&\cos{\theta^{\prime}}\end{pmatrix}\mathbf{p}^{\prime\rm T}+\mathbf{t}^{\prime}.$ (15)
For 2D minutiae, we also conduct pretraining for 80 epochs, followed by 100 epochs of finetuning. The other experimental conditions are the same as the last row in Table VII. It can be observed that the performance of 2D minutiae graph embedding is significantly worse than that of 3D minutiae graph embedding in both matching and identification. The 3D minutiae graph embedding can decrease the EER to 9.16%, and improve the Rank-1 accuracy to 88.35%, respectively. The improvement indicates that the graph embedding of 2D minutiae does not provide additional discriminative information for fingerprint recognition and even harms the discriminative ability of the original method. The improvement further confirms that our method introduces three-dimensional information into contactless fingerprint recognition, thereby enhancing the performance of contactless fingerprint recognition.
We further test the direct application of the pre-trained model for matching and identification experiments, which is listed as the ’2.5D Minutiae’ in Table VIII. As described in Section III-C3, the pre-trained model is only trained on the expansion dataset from contact-based 2D fingerprints to simulated 3D fingerprints. These fingerprints are not real 3D fingerprints. Therefore, they are listed as ’2.5D Minutiae’ in Table VIII. The remaining experimental settings are the same as the 3D minutiae graph embedding network.
The matching and identification performances are also shown in TABLE VIII and Fig. 13. The results obtained using only the pre-trained model are very close to those of 2D minutiae. Although the minutiae trained by the pretrained model are still in 3D, all the minutiae share the same 3D surface. Therefore,the pretrained model also does not introduce new three-dimensional information. However, our model fails to converge on the augmented 3D minutiae data without pretraining. Therefore, pretraining is necessary.
TABLE VIII: The comparison results between 2D minutiae and 3D minutiae on the UWA database
EER Rank-1
2D Minutiae
19.22% 87.95%
2.5D Minutiae
19.02% 88.00%
3D Minutiae
9.16% 88.35%
Figure 9: DET, ROC and CMC curves of the UWA database following matching protocols in Cui et al. [18]
Figure 10: DET, ROC and CMC curves of the UWA database following matching protocols in Dong et al. [17]
Figure 11: DET, ROC and CMC curves of the CFPose database following matching protocols in Dong et al. [17]
Figure 12: DET, ROC and CMC curves of ablation study of training strategies on the UWA database
Figure 13: DET, ROC and CMC curves of ablation study of 2D minutiae and 3D minutiae methods on the UWA database
V Conclusion
In this paper, we propose a novel contactless fingerprint recognition method that uses 3D graph matching to implement a 3D matching algorithm for contactless fingerprints. In contrast, previous methods only match fingerprints in 2D space and neglect the 3D information during matching. Experiment results on two public datasets prove that our method reaches state-of-the-art matching performances, and can successfully match contactless fingerprints across different poses.
Acknowledgments
We would like to appreciate Mr. Hanzhuo Tan for providing us with the CFPose database for experiments.
This work is supported in part by the National Natural Science Foundation of China under Grants 62206026.

### v2

@@ Appendix B Supplement to the Feature Extraction Experiments
VeriFinger
51.7% 74.9% 54.9%
Tan and Kumar[tan2020towards]
65.4% 70.1% 67.2%
Proposed
68.7% 67.1% 67.4%
Table 4: Minutiae extraction at 8-pixel threshold results
Appendix C Supplement to the Matching Experiments
We conduct experiments on UWA separately according to Cui et al.’s [cui2023monocular] protocol which performs $\times3,000$ matchings. There are 9,000 pairs of genuine matchings, and the rest of the pairs are impostor matching. The missing partial matching scores are set to zero. For evaluating the matching performance on CFPose, a total of 719,400 pairs of matchings are performed. 5,400 pairs are genuine matching, and the rest are impostor matching. As segmentation has minimal impact on fingerprint matching, we still use all 9,888 fingerprints for matching experiment. Following [dong2023syn], we utilize $824\times 12\times 11/2=54,384$ genuine matches and $824\times 823/2=339,076$ impostor matches (from the first image of each id) for matching experiments.
Fig. 10 presents the DET curves of our method on the CFPose and ZJU datasets. It shows that our new depth prediction and graph matching method significantly boosts contactless fingerprint recognition, especially when unfolding CFPose for 2D matching. This improves fingerprint matching performance and addresses previous poor generalization in fingerprint depth prediction.
Table 5 shows the ZeroFMR of different methods. Upon combining with the DET curve we provide, it can be found that our method surpasses previous ones at various FMR levels.
Method ZeroFMR
CFPose[tan2020towards] UWA[zhou2014benchmark] ZJU[grosz2022contact]
Verifinger 7.04% 66.64% 1.66%
Tan and Kumar [tan2020towards] 5.43% 44.53% -
Cui et al. [cui2023monocular] - 51.99% -
Proposed 3.37% 44.94% 1.39%
Proposed 3.65% 43.23% 1.16%
  • $\dagger$}: Only unwarp by our predicted depth for 2D matching.
Table 5: ZeroFMR on three databases
Appendix D Ablation Study
a b c d e f EER
21.99%
21.46%
18.99%
21.00%
16.80%
11.16%
9.23%
  • a: Minutiae from Verifinger.
  • b: Minutiae from our 3D Feature Network.
  • c: Combining 128-dim deep feature for training.
  • d: Only using 2D Minutiae
@@ Appendix D Ablation Study
  • f: 2D Score with our unwarp method.
Table 6: Ablation study of training strategies on the UWA database
Table 6 presents the results of our ablation experiments. This ablation study focuses on the 3D matching process. First, by replacing the minutiae extracted by Verifinger with those from our 3D Feature Network, we demonstrate that our network better captures the shape of 3D fingerprints. Next, incorporating the 128-dimensional deep features from the intermediate layer of the 3D Feature Network further enhances network performance. We also show that position correction in 3D space improves performance. Finally, our method’s ability to effectively combine with other approaches significantly boosts the performance of cross-pose contactless fingerprint recognition. At the same time, our method can also capture new matching information on the ZJU dataset, which lacks large scale pose changes. It works even when FMR is high, effectively handling difficult samples.
Moreover we re-trained the 3D Graph Matching Network using 2D minutiae and evaluated it directly on the UWA dataset for matching. The 3D minutiae results below correspond to the 3D matching outcomes in our Supplementary Material when no 2D scores were fused. To align the input formats of 2D and 3D minutiae, instead of using the traditional $(x,y,\theta)$ as input, we used $(x,y,d_{x},d_{y})$. The experimental result shows that 3D minutiae can significantly enhance the performance of 3D graph matching.